# Topological characterization of phase transitions and critical edge states in one-dimensional non-Hermitian systems with sublattice symmetry

### v1

@@ I Introduction
Recently, the study of symmetry-protect topological (SPT) phases has been extended to gapless situations [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84]. The theoretical description of a gapless SPT (gSPT) state not only goes beyond the Landau paradigm of phase transitions based on local order parameters and symmetry-breaking [85], but also falls outside the standard classification of topological insulators and superconductors following the tenfold way [86]. In the simplest case, a gSPT phase can be realized by a critical point between two topologically nontrivial phases [69]. In one spatial dimension, such a realization could be the transition point between two gapped phases with different and nonzero topological invariants in a spin or superconducting chain [47], and the nontrivial topology of the critical point is signified by the presence of degenerate edge modes at the gap-closing point of the bulk energy spectrum [54]. Be coexisting with a gapless bulk, these topological edge modes can also be regarded as critical. In the presence of non-Hermitian effects, the description of gSPT phases could become more complicated. On the one hand, unique non-Hermitian topology related to the exceptional point and non-Bloch band theory may lead to new forms of gSPT phases beyond closed-system constraints. On the other hand, gain and loss or nonreciprocal effects may influence the existence and stability of critical edge modes in an originally Hermitian gSPT phase. Addressing these issues could not only extend the conception of gSPT phases to non-Hermitian setups, but also offer potential insights for further explorations of mixed-state topology [87, 88, 89, 90, 91, 92] and bulk-edge correspondence in gapless open systems.
@@
In this work, we propose a theoretical framework to characterize topologically nontrivial critical points and critical edge states in one-dimensional (1D), two-band non-Hermitian systems. Our approach is applicable to the description of both gapped and gapless sublattice-symmetry-protected topological phases. Moreover, it does not concern whether the underlying non-Hermitian band theory is in Bloch (with standard Brillouin zone) or non-Bloch (with generalized Brillouin zone) form. The rest of the paper is organized as follow. In Sec. II, we formulate the definition of topological invariants and the related bulk-edge correspondence in our theory. The topological invariants are deduced from the algebraic zero-pole counting of two characteristic polynomials for a given non-Hermitian Hamiltonian. In Sec. III, we apply our theory to characterize the topological phases, phase transitions and edge states in a broad class of non-Hermitian Su-Schrieffer-Heeger (NHSSH) chains, with a focus on the topological nature and bulk-edge correspondence at their gapless critical points. Different classes of topologically distinct critical points and phase transitions along topological phase boundaries are revealed, and the configuration of their critical edge modes are identified. In Sec. IV, we summarize our results and discuss potential future studies. Some calculation details about the non-Hermitian topological edge states are presented in Appendix A.
@@ II Theory
In this section, we outline the theoretical framework we proposed to characterize gapped and gapless topological phases in 1D, two-band non-Hermitian systems with sublattice symmetry. Despite showing the basic formulation, we also discuss the issues of some previous theories on the topological characterization of phase transition points with vanishing bulk gaps and critical edge states. The application of our theory to concrete models is presented in the next section.
@@
where $k\in[-\pi,\pi]$ is the quasimomentum. The off-diagonal elements $f(k)$ and $g(k)$ are both $2\pi$-periodic in $k$, and the $H(k)$ is non-Hermitian if there exists a quasimomentum $k$ such that $f(k)\neq g^{*}(k)$. The sublattice symmetry of $H(k)$ is given by $\Gamma=\sigma_{z}$, in the sense that $\Gamma H(k)\Gamma=-H(k)$ and $\Gamma^{2}=\sigma_{0}$, where $\sigma_{0}$ denotes the $2\times 2$ identity matrix. The Hamiltonian $H(k)$ has two energy bands, whose dispersion relations are given by
$E_{\pm}(k)\equiv\pm E(k)=\pm\sqrt{f(k)g(k)}.$ (2)
$\frac{b(k)}{a(k)}=\frac{E(k)}{f(k)}=\frac{g(k)}{E(k)}.$ (3)
$P(k)\equiv\frac{g(k)}{f(k)},\qquad Q(k)\equiv f(k)g(k).$ (4)
As the $f(k)$ and $g(k)$ are $2\pi$-periodic in $k$, the characteristic functions $P(k)$ and $Q(k)$ should also satisfy $P(k+2\pi)=P(k)$ and $Q(k+2\pi)=Q(k)$. We could thus expand the $P(k)$ and $Q(k)$ into Fourier series as $P(k)=\sum_{n\in\mathbb{Z}}p_{n}e^{ink}$ and $Q(k)=\sum_{n\in\mathbb{Z}}q_{n}e^{ink}$. These series can be viewed as polynomials of the exponential factor $e^{ik}$. To treat Bloch and non-Bloch band theories of non-Hermitian systems on an equal footing, we extend the $e^{ik}$ to the whole complex plane by setting $e^{ik}\rightarrow z$ with $z\in\mathbb{C}$. The resulting characteristic polynomials of the complex-extended Hamiltonian $H(z)$ are given by $P(z)=\sum_{n\in\mathbb{Z}}p_{n}z^{n}$ and $Q(z)=\sum_{n\in\mathbb{Z}}q_{n}z^{n}$, or
@@
where $f(z)\equiv f(k\rightarrow-i\ln z)$ and $g(z)\equiv g(k\rightarrow-i\ln z)$ are “complex continuations” of $f(k)$ and $g(k)$. Inside a closed contour on the complex $z$-plane, the difference between the numbers of zeros and poles (including their multiplicities and orders) of $P(z)$ and $Q(z)$ yield two integer-quantized winding numbers $W_{1}$ and $W_{2}$, respectively, according to the Cauchy’s argument principle. Our theoretical proposal is based on the topological messages encoded in these winding numbers.
@@
In previous studies [27, 30], it was found that by identifying the first $2\mu=2\max\{m,n\}$ zeros and/or poles of $P(z)$ (ordered according to their magnitudes), we would obtain a winding number $W_{0}\equiv N_{z}^{\in 2\mu}-N_{p}^{\in 2\mu}$, which could characterize each topological phase of a 1D, two-band, sublattice-symmetric non-Hermitian system and determine the number $N_{0}$ of its edge zero modes via the relation $N_{0}=|W_{0}|$. Here, $m$ and $n$ are given by the highest negative powers of $f(z)$ and $g(z)$ in $z$. $N_{z}^{\in 2\mu}$ and $N_{p}^{\in 2\mu}$ are the numbers of zeros and poles of $P(z)$ inside the set of its first $2\mu$ zeros/poles. This algebraic approach is proved to be equivalent to other methods based on the generalized Brillouin zone (GBZ) and non-Bloch band theory for gapped phases [25, 26, 30]. However, this $P(z)$-based theory may lead to ambiguous predictions about the non-Hermitian topology and edge states at phase transition points, where the Bloch or non-Bloch bands of the system are touched at zero energy.
The above point can be made clear via a counterexample. Let us consider a complex-continued non-Hermitian Hamiltonian $H(z)$ with off-diagonal elements $f(z)=J_{0}+J_{1}z^{-1}$ and $g(z)=J_{0}+J_{1}z$. When $J_{0}$ and $J_{1}$ take real values, we go back to the standard SSH model after setting $z\rightarrow e^{ik}$. Following the $P(z)$-based approach, we have $m=1$, $n=0$ and thus $2\mu=2\max\{1,0\}=2$. By definition, we find $P(z)=g(z)/f(z)=z(J_{0}+J_{1}z)/(J_{0}z+J_{1})$. It can then be identified that if $|J_{1}|<|J_{0}|$, the first $2$ zeros/poles of $P(z)$ contains a zero at $z=0$ and a pole at $z=-J_{1}/J_{0}$, yielding $N_{z}^{\in 2}=1$, $N_{p}^{\in 2}=1$, $W_{0}=N_{z}^{\in 2\mu}-N_{p}^{\in 2\mu}=0$ and $N_{0}=|W_{0}|=0$. Meanwhile, if $|J_{1}|>|J_{0}|$, the first $2$ zeros/poles of $P(z)$ contains two zeros at $z=0$ and $z=-J_{0}/J_{1}$, yielding $N_{z}^{\in 2}=2$, $N_{p}^{\in 2}=0$, $W_{0}=N_{z}^{\in 2\mu}-N_{p}^{\in 2\mu}=2$ and $N_{0}=|W_{0}|=2$. These predictions are all correct, as the two bulk bands of $H(z)$ are gapped at zero energy $Q(z)\neq 0$] for any finite $z$ whenever $|J_{1}|\neq|J_{0}|$. However, at the phase transition point $|J_{1}|=|J_{0}|$, we have $|-J_{1}/J_{0}|=|-J_{0}/J_{1}|=1$, which means that we could not order the magnitudes of the zero and pole at $z=-J_{0}/J_{1}$ and $z=-J_{1}/J_{0}$. The determination of which are the first two zeros and/or poles of $P(z)$ then becomes ambiguous. If we include or exclude both their contributions to $W_{0}$, we will get $W_{0}=1$ and thus $N_{0}=1$. This is impossible, as the zero modes must come in pairs (i.e., $N_{0}\in 2\mathbb{N}$) due to the sublattice symmetry. If we only retain the zero at $z=-J_{0}/J_{1}$ as the second zero/pole, we will get $W_{0}=2$ and $N_{0}=2$, which is incorrect even for the Hermitian SSH model [46]. If we instead retain the pole at $z=-J_{1}/J_{0}$ as the second zero/pole, we will find $W_{0}=0$ and $N_{0}=0$. Even though these are the correct values of topological invariants and edge states at the critical point of standard SSH model [46], they could not be conclusively and unambiguously predicted by the $P(z)$-based approach. Therefore, the algebraic scheme in Refs. [27, 30] tends out to be inapplicable to the characterization of nontrivial topology and bulk-edge correspondence at the phase transition (bulk-gap closing) point of non-Hermitian systems. Similarly, as the algebraic and GBZ schemes are proved to be equivalent for sublattice-symmetric models [30], the standard GBZ approach is also expected to be unworkable for describing gSPT phases in non-Hermitian systems.
@@
These observations motivate us to generalize existing theories in order to characterize the nontrivial topology at non-Hermitian critical points. The above example suggests that there could be competing zeros and poles of equal magnitudes in $P(z)$, whose topological signatures may subject to cancellations at the critical point. Their contributions are instead all retained in the characteristic function $Q(z)$ as its zeros at the gap closing point. Therefore, we propose to introduce an additional winding number by applying the Cauchy’s argument principle to $Q(z)$. More precisely, we take the GBZ as the integration contour of the argument principle. The winding numbers of $P(z)$ and $Q(z)$ are then given by
@@
Here, $N_{1z}^{\in{\rm GBZ}}$ and $N_{1p}^{\in{\rm GBZ}}$ ($N_{2z}^{\in{\rm GBZ}}$ and $N_{2p}^{\in{\rm GBZ}}$) denote the numbers of zeros and poles of $P(z)$ ($Q(z)$) inside the GBZ. Note that the definition of $W_{1}$ is consistent with the definition of $W_{0}$ in the existing theory [30]. Combining the information encoded in $W_{1}$ and $W_{2}$, we arrive at the final topological invariant $W$ and bulk-edge correspondence of 1D, two-band systems with sublattice symmetry, i.e.,
@@
One may now check the applicability of our theory to the simplest SSH model. Besides the $P(z)=z(J_{0}+J_{1}z)/(J_{0}z+J_{1})$, we also have $Q(z)=(J_{0}z+J_{1})(J_{0}+J_{1}z)/z$. The GBZ here is just the standard Brillouin zone, which forms a circle of radius $1$ on the complex plane regardless of whether the values of $J_{0}$ and $J_{1}$ are real or complex. Direct calculations following Eqs. (6)–(8) predict $(W_{1},W_{2})=(0,0)$ when $|J_{1}|<|J_{0}|$, $(W_{1},W_{2})=(2,0)$ when $|J_{1}|>|J_{0}|$, and $(W_{1},W_{2})=(1,-1)$ when $|J_{1}|=|J_{0}|$, yielding $W=0,1,0$ and $N_{0}=0,2,0$ in these three parameter regions. Besides reproducing known results about gapped phases [93], we also confirm the topological triviality of the gapless phase boundary $|J_{1}|=|J_{0}|$ of the SSH model [47] without ambiguity. In the next section, we utilize our theory to depict the topology and edge states in both gapped phases and along gapless phase boundaries for a broad class of sublattice-symmetric non-Hermitian models.
@@ III Results
Figure 1: Schematic illustration of NHSSH $\alpha$ chains with $\alpha=0,1,2$. In each case, the solid and dotted curves with arrows denote hopping amplitudes from left to right and right to left between sublattices A (cyan balls) and B (orange balls).
In this section, we demonstrate the applicability of our theory to the characterization of topologically nontrivial critical points and zero-energy edge modes in non-Hermitian systems. We first introduce a general class of 1D non-Hermitian models with sublattice symmetry in Sec. III.1, which will be referred to as NHSSH $\alpha$ chains. Our theory is then applied to describing gapped topological phases, distinguishing between topologically trivial and nontrivial critical points, and depicting topological phase transitions without gap closing/reopening in Secs. III.2, III.3, and III.4, respectively, for three different types of NHSSH $\alpha$ chains. Theoretical analyses are complemented by numerical calculations of the spectra and edge states in each case.
@@ III.1 Model
The $\hat{a}_{n}^{\dagger}$ ($\hat{b}_{n}^{\dagger}$) creates a particle in the sublattice A (B) of unit cell $n$. $J_{\alpha}^{{\rm L}}$ ($J_{\alpha}^{{\rm R}}$) denotes the hopping amplitude from right to left (left to right) over $\alpha$ lattice sites. The $\hat{H}$ is non-Hermitian whenever there exists an $\alpha$ such that $J_{\alpha}^{{\rm L}}\neq(J_{\alpha}^{{\rm R}})^{*}$. The Hermitian SSH model is recovered by setting $\hat{H}=\sum_{\alpha=0,1}\hat{H}_{\alpha}$ and $J_{\alpha}^{{\rm L}}=J_{\alpha}^{{\rm R}}=J_{\alpha}\in\mathbb{R}$. The lattice geometries of $\hat{H}_{\alpha}$ with $\alpha=0,1,2$ are shown in Fig. 1.
@@
where $N$ is the total number of unit cells and $k\in[-\pi,\pi]$ denotes the quasimomentum. In momentum space, the component $\hat{H}_{\alpha}$ has the form $\hat{H}_{\alpha}=\sum_{k}\hat{\Psi}_{k}^{\dagger}H_{\alpha}(k)\hat{\Psi}_{k}$, where $\hat{\Psi}_{k}^{\dagger}\equiv(\hat{a}_{k}^{\dagger},\hat{b}_{k}^{\dagger})$. The $H_{\alpha}(k)$ is a $2\times 2$ off-diagonal matrix of the form
$H_{\alpha}(k)=\begin{bmatrix}0&J_{\alpha}^{{\rm R}}e^{-ik\alpha}\\ J_{\alpha}^{{\rm L}}e^{ik\alpha}&0\end{bmatrix}$. ($1$)
@@
The extended Bloch Hamiltonian of the NHSSH $\alpha$ chain is then given by $H(z)=\sum_{\alpha\in\mathbb{Z}}H_{\alpha}(z)$.
@@
In the following, we consider three different types of NHSSH $\alpha$ chains. We apply our theory to characterize the topological phases and bulk-edge correspondence of the system in each case.
III.2 Single $\alpha$ chain: gapped and gapless phases
The NHSSH single $\alpha$ chain has the complex-extended Hamiltonian $H(z)=H_{\alpha}(z)$. It contains only a single length scale in its hopping amplitudes $J_{\alpha}^{{\rm L,R}}$. In this subsection, we assume $\alpha\geq 0$ without losing generality.
$P(z)=\frac{J_{\alpha}^{{\rm L}}}{J_{\alpha}^{{\rm R}}}z^{2\alpha},\quad Q(z)=J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}.$ (13)
Since the $Q(z)=[E(z)]^{2}$ is independent of $z$, the bulk spectrum of the system is formed by two flat bands at $\pm E(z)=\pm\sqrt{J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}}$, and the GBZ is just the standard BZ with $k\in[-\pi,\pi]$. On the complex $z$-plane, the $Q(z)$ has neither zeros nor poles, while the $P(z)$ has a zero of order $2\alpha$ at $z=0$, which is inside the BZ unit circle $|z|=1$. According to the definition of winding numbers for $P(z)$ and $Q(z)$ in Eq. (6), we find $W_{1}=2\alpha$ and $W_{2}=0$. The topological invariant $W$ of the system and the number of its degenerate zero modes $N_{0}$ under the open boundary condition (OBC) are thus given by Eqs. (7) and (8) as
$W=\frac{W_{1}+W_{2}}{2}=\alpha,\quad N_{0}=2|W|=2\alpha.$ (14)
We now verify our results about the NHSSH single $\alpha$ chain with two groups of numerical examples. In the first group, we consider the case of asymmetric hopping with $(J_{\alpha}^{{\rm L}},J_{\alpha}^{{\rm R}})=(0.5,2)$ for $\alpha=0,1,2$. The bulk spectrum is expected to show two flat bands at $\pm\sqrt{J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}}=\pm 1$, separated by a constant gap $\Delta E=2$. In the second group, we consider the case of unidirectional hopping with $(J_{\alpha}^{{\rm L}},J_{\alpha}^{{\rm R}})=(0,1)$ for $\alpha=0,1,2$. The bulk spectrum is expected to show two degenerate flat bands at zero energy. In these three cases, the theoretically predicted winding numbers and numbers of zero-energy edge modes are $W=(0,1,2)$ and $N_{0}=(0,2,4)$ according to Eq. (14), regardless of the chosen parameter values $(J_{\alpha}^{{\rm L}},J_{\alpha}^{{\rm R}})$.
Figure 2: Spectra and edge states of NHSSH single $\alpha$ chains with bulk gaps. The system has $10$ unit cells and 2020 lattice sites. $j$ is the state index. $l$ is the lattice index. (a) Energy spectrum of $\hat{H}_{0}$ with $(J_{0}^{{\rm L}},J_{0}^{{\rm R}})=(0.5,2)$ under the PBC (in circles) and OBC (in stars). (b) Energy spectrum of $\hat{H}_{1}$ with $(J_{1}^{{\rm L}},J_{1}^{{\rm R}})=(0.5,2)$ under the PBC (in circles) and OBC (in stars). (c) Energy spectrum of $\hat{H}_{2}$ with $(J_{2}^{{\rm L}},J_{2}^{{\rm R}})=(0.5,2)$ under the PBC (in circles) and OBC (in stars). (d) Probability distributions $P$ of the zero-energy edge states of $\hat{H}_{1}$ in (b). (e) Probability distributions $P$ of the zero-energy edge states of $\hat{H}_{2}$ in (c). In (d) and (e), different edge states are distinguished by the colors of probability bars.
Figure 3: Spectra and edge states of NHSSH single $\alpha$ chains without bulk gaps. The system has $10$ unit cells and 2020 lattice sites. $j$ is the state index. $l$ is the lattice index. (a) Energy spectrum of $\hat{H}_{0}$ with $(J_{0}^{{\rm L}},J_{0}^{{\rm R}})=(0,1)$ under the PBC (in circles) and OBC (in stars). (b) Energy spectrum of $\hat{H}_{1}$ with $(J_{1}^{{\rm L}},J_{1}^{{\rm R}})=(0,1)$ under the PBC (in circles) and OBC (in stars). (c) Energy spectrum of $\hat{H}_{2}$ with $(J_{2}^{{\rm L}},J_{2}^{{\rm R}})=(0,1)$ under the PBC (in circles) and OBC (in stars). (d) Probability distributions $P$ of the zero-energy edge states of $\hat{H}_{1}$ in (b). (e) Probability distributions $P$ of the zero-energy edge states of $\hat{H}_{2}$ in (c). In (d) and (e), different edge states are distinguished by the colors of probability bars.
In Fig. 3, we present the energy spectra and edge states of the single $\alpha$ chain for $\alpha=0,1,2$ and $(J_{\alpha}^{{\rm L}},J_{\alpha}^{{\rm R}})=(0,1)$. The spectrum of $\hat{H}_{0}$ in Fig. 3(a) has no edge zero modes ($N_{0}=0$). The spectrum of $\hat{H}_{1}$ in Fig. 3(b) holds two zero-energy edge modes ($N_{0}=2$), whose distributions are shown in Fig. 3(d). The spectrum of $\hat{H}_{2}$ in Fig. 3(c) contains four zero-energy edge modes ($N_{0}=4$), whose distributions are shown in Fig. 3(e). These results again verify our theoretical predictions. Notably, even though the bulk spectra of $\hat{H}_{0,1,2}$ have no gaps at $E=0$ in Figs. 3(a)–3(c), our theory could still correctly predict the numbers of edge zero modes that can coexist with gapless bulks in these cases.
The presence or absence of edge zero modes in the spectrum of $\hat{H}_{\alpha}$ for a given $\alpha$ can be understood from the geometric connectivity of the underlying lattice (see Fig. 1). For $\hat{H}_{0}$, the sublattices A and B in each unit cell are coupled due to the non-vanishing $J_{0}^{{\rm L}}$ and/or $J_{0}^{{\rm R}}$, making it impossible to have isolated sites at boundaries to host edge zero modes. For $\hat{H}_{1}$, the sublattice B in unit cell $n$ is coupled to the sublattice A in unit cell $n+1$ for n=1,,N1n=1,...,N-1. The sublattices A in unit cell $1$ and B in unit cell $N$ are thus isolated from other sites, making them available to host two edge zero modes $\hat{a}_{1}^{\dagger}|\emptyset\rangle$ and $\hat{b}_{N}^{\dagger}|\emptyset\rangle$, where $|\emptyset\rangle$ denotes the vacuum state. For $\hat{H}_{2}$, the sublattice B in unit cell $n$ is coupled to the sublattice A in unit cell $n+2$ for n=1,,N2n=1,...,N-2. The sublattices A in unit cells $1,2$ and B in unit cells N1,NN-1,N are thus isolated from other sites, making them available to host four edge zero modes $\hat{a}_{1}^{\dagger}|\emptyset\rangle$, $\hat{a}_{2}^{\dagger}|\emptyset\rangle$, $\hat{b}_{N-1}^{\dagger}|\emptyset\rangle$ and $\hat{b}_{N}^{\dagger}|\emptyset\rangle$. It is also clear that the topological origin of zero-energy edge modes in a single $\alpha$ chain does not depend on whether the system is Hermitian or non-Hermitian.
In the next two subsections, we study NHSSH $\alpha$ chains with topologically distinct critical points and critical edge zero modes.
@@ III.2 Single $\alpha$ chain: gapped and gapless phases
III.3 $(\alpha,\alpha^{\prime})$ chain: topologically trivial/nontrivial critical points

### v2

@@ I Introduction
Recently, the study of symmetry-protect topological (SPT) phases has been extended to gapless situations [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88]. The theoretical description of a gapless SPT (gSPT) state not only goes beyond the Landau paradigm of phase transitions based on local order parameters and symmetry-breaking [89], but also falls outside the standard classification of topological insulators and superconductors following the tenfold way [90]. In the simplest case, a gSPT phase can be realized by a critical point between two topologically nontrivial phases [72]. In one spatial dimension, such a realization could be the transition point between two gapped phases with different and nonzero topological invariants in a spin or superconducting chain [47], and the nontrivial topology of the critical point is signified by the presence of degenerate edge modes at the gap-closing point of the bulk energy spectrum [54]. Be coexisting with a gapless bulk, these topological edge modes can also be regarded as critical. In the presence of non-Hermitian effects, the description of gSPT phases could become more complicated. On the one hand, unique non-Hermitian topology related to the exceptional point and non-Bloch band theory may lead to new forms of gSPT phases beyond closed-system constraints. On the other hand, gain and loss or nonreciprocal effects may influence the existence and stability of critical edge modes in an originally Hermitian gSPT phase. Addressing these issues could not only extend the scope of gSPT phases to non-Hermitian setups, but also offer potential insights for further explorations of mixed-state topology [91, 92, 93, 94, 95, 96] and bulk-edge correspondence in gapless open systems.
Figure 1: Schematic diagrams of the theory and model. (a) and (b) illustrate the configurations of zeros (in purple dots) and poles (in green circles) of the characteristic functions $P(z)$ and $Q(z)$ [Eq. (2)] on the complex $z$ plane, with the GBZ given by the irregularly shaped contours. Only zeros/poles inside the GBZ contribute to the winding numbers $W_{1}$ and $W_{2}$ [Eq. (3)], whose combination predicts the topological invariant $W=1$ [Eq. (4)] and number of edge zero modes $N_{0}=2$ [Eq. (5)] for the case illustrated in (a)–(b). (c) shows NHSSH $\alpha$ chains with $\alpha=0,1,2$. In each case, the solid and dotted curves with arrows denote hopping amplitudes from left to right and right to left between sublattices A (cyan balls) and B (orange balls).
@@
In this work, we propose a theoretical framework to characterize topologically nontrivial critical points and critical edge states in one-dimensional (1D), two-band non-Hermitian systems. Our approach is applicable to the description of both gapped and gapless sublattice-symmetry-protected topological phases. Moreover, it does not concern whether the underlying non-Hermitian band theory is in Bloch (with standard Brillouin zone) or non-Bloch (with generalized Brillouin zone) form. The rest of the paper is organized as follow. In Sec. II, we formulate the definition of topological invariants and the related bulk-edge correspondence in our theory. The topological invariants are deduced from the algebraic zero-pole counting of two characteristic polynomials for a given non-Hermitian Hamiltonian [Figs. 1(a)–1(b)]. In Sec. III, we apply our theory to characterize the topological phases, phase transitions and edge states in a broad class of non-Hermitian Su-Schrieffer-Heeger (NHSSH) chains, with a focus on the topological nature and bulk-edge correspondence at their gapless critical points. Different classes of topologically distinct critical points and phase transitions along topological phase boundaries are revealed, and the configuration of their critical edge modes are identified. In Sec. IV, we summarize our results and discuss potential future studies. Some further theoretical details and model illustrations are presented in Appendices AC.
@@ II Theory
In this section, we outline the theoretical framework we proposed to characterize gapped and gapless topological phases in 1D, two-band non-Hermitian systems with sublattice symmetry. Despite showing the basic formulation, we also discuss the issue of some previous theories on the topological characterization of phase transition points with vanishing bulk gaps and critical edge states. The application of our theory to concrete models is presented in the next section. Some further theoretical details are provided in Appendix A.
@@
where $k\in[-\pi,\pi]$ is the quasimomentum. The $H(k)$ in Eq. (1) possesses the sublattice symmetry $\Gamma=\sigma_{z}$. Its phases at half-filling are then characterized by integer-valued topological numbers. To consistently define these invariants for both gapped and gapless phases, we introduce a pair of characteristic functions for the complex-extended Hamiltonian $H(z)\equiv H(k\rightarrow-i\ln{z})$, which are given by [see Appendix A for further details and Figs. 1(a)–1(b) for graphic illustrations]
@@
where $f(z)\equiv f(k\rightarrow-i\ln z)$ and $g(z)\equiv g(k\rightarrow-i\ln z)$ are “complex continuations” of the $f(k)$ and $g(k)$ in $H(k)$ to the whole complex plane $z\in\mathbb{C}$. Inside a closed contour on the complex plane, the difference between the numbers of zeros and poles (including their multiplicities and orders) of $P(z)$ and $Q(z)$ yield two integer winding numbers $W_{1}$ and $W_{2}$, respectively, according to the Cauchy’s argument principle. Our theoretical proposal is based on the topological messages encoded in these winding numbers.
@@
In previous studies [27, 30], it was found that by identifying the first $2\mu=2\max\{m,n\}$ zeros and/or poles of $P(z)$ (ordered according to their magnitudes), we would obtain a winding number $W_{0}\equiv N_{z}^{\in 2\mu}-N_{p}^{\in 2\mu}$, which could characterize each topological phase of a 1D, two-band, sublattice-symmetric non-Hermitian system and determine the number $N_{0}$ of its edge zero modes via the relation $N_{0}=|W_{0}|$. Here, $m$ and $n$ are given by the highest negative powers of $f(z)$ and $g(z)$ in $z$. $N_{z}^{\in 2\mu}$ and $N_{p}^{\in 2\mu}$ are the numbers of zeros and poles of $P(z)$ inside the set of its first $2\mu$ zeros/poles. This algebraic approach is proved to be equivalent to other methods based on the generalized Brillouin zone (GBZ) and non-Bloch band theory for gapped phases [25, 26, 30]. However, this $P(z)$-based theory may lead to ambiguous predictions about the non-Hermitian topology and edge states at phase transition points, where the Bloch or non-Bloch bands of the system are touched at zero energy. We discuss an explicit example to showcase this ambiguity in Appendix A.
@@
At critical points, a key issue of the existing theory is that there could be competing zeros and poles of equal magnitudes in $f(z)$ and $g(z)$, whose topological signatures may subject to cancellations in $P(z)$ by definition. Their contributions are instead retained in the function $Q(z)$ as its zeros at the gap closing point. Therefore, we propose to introduce an additional winding number by applying the Cauchy’s argument principle to $Q(z)$. More precisely, by taking GBZ as the integration contour, the winding numbers of $P(z)$ and $Q(z)$ are given by
@@
Here, $N_{1z}^{\in{\rm GBZ}}$ and $N_{1p}^{\in{\rm GBZ}}$ ($N_{2z}^{\in{\rm GBZ}}$ and $N_{2p}^{\in{\rm GBZ}}$) denote the numbers of zeros and poles of $P(z)$ ($Q(z)$) inside the GBZ. Note that the definition of $W_{1}$ here is consistent with that of $W_{0}$ in the existing theory [30]. Combining the information of $W_{1}$ and $W_{2}$, we arrive at the final topological invariant $W$ and bulk-edge correspondence of 1D, two-band systems with sublattice symmetry, i.e.,
@@
In Appendix A, we illustrate the applicability of our theory to the simplest SSH model away from and at its criticality. In the next section, we utilize our theory to depict the topology and edge states in both gapped phases and along gapless phase boundaries for a broad class of sublattice-symmetric non-Hermitian models.
@@ III Results
In this section, we demonstrate the applicability of our theory to the characterization of topologically nontrivial critical points and zero-energy edge modes in non-Hermitian systems. We first introduce a general class of 1D non-Hermitian models with sublattice symmetry in Sec. III.1, which will be referred to as NHSSH $\alpha$ chains. Our theory is then applied to describing gapped topological phases, distinguishing between topologically trivial and nontrivial critical points, and depicting topological phase transitions without gap closing/reopening in Secs. III.2 and III.3. Theoretical analyses are complemented by numerical calculations of the spectra and edge states in each case. For completeness, we also discuss in Appendix B the topological phases for the simplest class of NHSSH $\alpha$ chains introduced in Sec. III.1, which could be either fully gapped or gapless without having topological phase transitions.
@@ III.1 Model
The $\hat{a}_{n}^{\dagger}$ ($\hat{b}_{n}^{\dagger}$) creates a particle in the sublattice A (B) of unit cell $n$. $J_{\alpha}^{{\rm L}}$ ($J_{\alpha}^{{\rm R}}$) denotes the directional hopping amplitude from sublattices A to B (B to A) over $\alpha$ lattice sites. The $\hat{H}$ is non-Hermitian whenever there exists an $\alpha$ such that $J_{\alpha}^{{\rm L}}\neq(J_{\alpha}^{{\rm R}})^{*}$. The Hermitian SSH model is recovered by setting $\hat{H}=\sum_{\alpha=0,1}\hat{H}_{\alpha}$ and $J_{\alpha}^{{\rm L}}=J_{\alpha}^{{\rm R}}=J_{\alpha}\in\mathbb{R}$. The lattice geometries of $\hat{H}_{\alpha}$ with $\alpha=0,1,2$ are shown in Fig. 1(c).
@@
where $N$ is the total number of unit cells and $k\in[-\pi,\pi]$. In momentum space, the component $\hat{H}_{\alpha}$ has the form $\hat{H}_{\alpha}=\sum_{k}\hat{\Psi}_{k}^{\dagger}H_{\alpha}(k)\hat{\Psi}_{k}$, where $\hat{\Psi}_{k}^{\dagger}\equiv(\hat{a}_{k}^{\dagger},\hat{b}_{k}^{\dagger})$. The $H_{\alpha}(k)$ is a $2\times 2$ off-diagonal matrix of the form
$H_{\alpha}(k)=\begin{bmatrix}0&J_{\alpha}^{{\rm R}}e^{-ik\alpha}\\ J_{\alpha}^{{\rm L}}e^{ik\alpha}&0\end{bmatrix}$. (8)
@@
The extended Bloch Hamiltonian of the NHSSH $\alpha$ chain is then given by $H(z)=\sum_{\alpha\in\mathbb{Z}}H_{\alpha}(z)$. The topological properties of $H_{\alpha}(z)$ with a fixed $\alpha$ is discussed in Appendix B.
@@
In the following, we consider two different types of NHSSH $\alpha$ chains, and apply our theory to characterize the topological phases and bulk-edge correspondence of these systems.
@@
III.2 $(\alpha,\alpha^{\prime})$ chain: topologically trivial/nontrivial critical points